OpenStudy (anonymous):

selling price of 30 articles is equal to the cost price of 20 articles. find the loss or profit percentage? please answer quickly

OpenStudy (anonymous):

Let = selling price = cost price. The article cost \[\frac{ P }{ 20}\]dollars each. they sold for \[\frac{ P }{ 30 }\] dollars each.

OpenStudy (anonymous):

You do \[\frac{ P }{ 30 } - \frac{ P }{ 20 } \]

OpenStudy (anonymous):

do you know how to solve from there ? :)